Springfield Corporation purchases a new machine on March 3, 20X4 for $35,600 in cash. It pays an additional $3,400 to transport and set up the machine. Springfield’s accountant determines that the equipment has no residual value and that the useful life is five years. It is expected to generate 2,400,000 units during its life. Assume Springfield employs the half-year convention. 

a. Record the purchase of the machine.

b. Assume that Springfield uses the straight-line method of depreciation. Record depreciation expense for the first two years of the machine’s life. 

c. Assume that Springfield uses the double-declining balance method of depreciation. Record depreciation expense for the first two years of the machine’s life. 

d. Assume that Springfield uses the units-of-production method of depreciation. 

During Year 1, the machine produces 600,000 units. 

During Year 2, the machine produces 578,000 units. 

Record depreciation expense for the first two years of the machine’s life


1
Expert's answer
2021-12-05T18:57:00-0500

a. Record the "(\$39,000 −\$7,800) × \frac{2}{5 }= \$12,480" purchase of the machine.

Date Account Title Debit Credit

Mar 3 20x4 Machine 39000

Cash 39000


b. Depreciation expense for the first two years of the machine’s life.


Depreciation Expense = "\frac{\$39,000}{5 years}"


Depreciation Expense = "\$7,800" per year

20X4: Take "\frac{1}{2}" of first year due to half year convention convention = "\frac{1}{2}\times\$7800=\$3900"


20X5: expense = "\$7,800"


c. Record depreciation expense for the first two years of the machine’s life.


Straight line rate = "\frac{1}{5}" Double declining rate = "\frac{1}{2}"


20X4 Depreciation Expense ="\$39,000 ×\frac{ 2}{5} × \frac{1}{2} = \$7,800"


20X5 Depreciation Expense ="(\$39,000 − \$7,800) × \frac{2}{5 }= \$12,480"


d. Assume that Springfield uses the units-of-production method of depreciation. Record depreciation expense for the first two years of the machine’s life.

Depreciation Rate ="\frac{ \$39,000}{2,400,000}" units


Depreciation Rate = 0.01625 per unit


20X4 Depreciation Expense ="600,000 units \times 0.0162"


20X4 Depreciation Expense = "\$9,750" (no need to adjust for half year since expense is based on ased on usage rather than time) usage rather than time)


20X5 Depreciation Expense = "578,000 units \times 0.01625"


20X5 Depreciation Expense = "\$9,393"